OpenStudy (anonymous):

The slope of a line is 9 and the point (9, 15) is on the line. Express the line in slope-intercept form.

OpenStudy (anonymous):

u should try then give me your answer what do you think about that?

OpenStudy (anonymous):

i got y = 15x + -144

OpenStudy (anonymous):

y-15 =9 ----- x-9 y-15=9(x-9) y-15=9x-81 y=9x-66 it cant be 15x since your gradient is 9 therefore its 9x

OpenStudy (anonymous):

what about this The slope of a line is 13 and the point (3, 19) is on the line. Express the line in slope-intercept form.

OpenStudy (anonymous):

you already know that x is 13 thus 13x y-19 = 13 ---- x-3 y-19=13(x-3) y-19=13x-39 y=13x-20